OpenStudy (anonymous):

ACT Question: If the value, to the nearest thousandth, of cos theta is -0.385, which of the following could be true about theta? a.) theta is greater than or equal to 0 and less than pi/6 b.) theta is greater than or equal to pi/6 and less than pi/3 c.) theta is greater than or equal to pi/3 and less than pi/2 d.) theta is greater than or equal to pi/2 and less than 2pi/3 e.) theta is greater than or equal to 2pi/3 and less than or equal to pi

OpenStudy (anonymous):

If the cosine of an angle is negative, then the angle is in either the second or third quadrant.

OpenStudy (anonymous):

@jim_thompson5910 @Jaynator495 @JoKeR0331 @jazzyspazzy @jennyrlz @Jess998776 @xX_GamerGirl_Xx @Xman1295 @Xaze @Lady.Liv1776 @LazyBoy @lily1021 @leonelaflores @leonelaflores @Lilthum @Lilthum @woodrow73 @esam2 @rational @TheSmartOne @yoitsjessie @iheartfood @pooja195 @adajiamcneal @SithsAndGiggles @dan815 @FloppyRawrRawr @hhelpplzzzz @Lady.Liv1776 @KyanTheDoodle @ZaneMarcs @coolman500 @vicyoria15 @nikato @BlossomCake @MTALHAHASSAN2

pooja195 (pooja195):

Mass tag much? There is already someone helping.

OpenStudy (anonymous):

Please explain it cause I know the answer from the back of the act book but I just dont get it

pooja195 (pooja195):

Im not helping. @SithsAndGiggles is already helping you why not try to understand them rather than bring attention to the post.

OpenStudy (anonymous):

In terms of \(x,y,z\), you have \(\cos\theta=\dfrac{x}{z}\) (where \(\theta\) is the angle I mistakenly left out). \(z\) is always taken to be a positive number, so if the trig ratio is negative, either \(x\) or \(y\) must be negative.

OpenStudy (anonymous):